OpenStudy (anonymous):

Still stuck on this! :( On the top of a hill stands a tower that is 50 feet tall. From a point that is on the same level as the base of the hill, the angle of elevation to the top of the tower is 56° and the angle to the bottom of the tower is 51°. What is the height of the hill?

OpenStudy (anonymous):

a. 37 b. 58 c. 249 d. 347

OpenStudy (anonymous):

i think i did solve this and got 250 already :P

OpenStudy (anonymous):

Oh I think you did! :/

OpenStudy (dls):

Don't practice double posting.

OpenStudy (anonymous):

@DLS I would not have "double posted" this problem if I had remembered that sami had already helped me on it.

OpenStudy (anonymous):

\[ \let the distance \between the measuring point and the hill="Y" andlet the of the hell="x" ;then at first \tan "56"=(X+5o)/Y ;then (1.5Y=50+X) at second \tan"51"=X/y then 1.2Y=x fromthe 2 equations"X"=249.3\approx249\]

OpenStudy (anonymous):

Hmm, your reply isn't showing up @hamidic :(

OpenStudy (anonymous):

its 249.3 im sure

OpenStudy (anonymous):

Thank you for your helpful input :) I appreciate it